I have asked this question before and it was wrong the answer so still don't understand. If someone could please help me?

  1. The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

    ACCOUNT Work in Process—Baking Department ACCOUNT NO.
    Date Item Debit Credit Balance
    Debit Credit
    March 1 Bal., 6,000 units, 3/5 completed 22,800
    31 Direct materials, 108,000 units 205,200 228,000
    31 Direct labor 55,600 283,600
    31 Factory overhead 31,280 314,880
    31 Goods finished, 109,500 units 304,170 10,710
    31 Bal. ? units, 3/5 completed 10,710

    a. Determine the number of units in work in process inventory at March 31.
    units

    Feedback

    How many units were in process during the period? How many units were completed during the period?

    b. Determine the equivalent units of production for direct materials and conversion costs in March. If an amount is zero, enter in "0".

    Baking Department
    Equivalent Units of Production for Direct Materials and Conversion Costs
    For March
    Whole Units Direct Materials
    Equivalent Units
    Conversion Equivalent Units
    Inventory in process, March 1
    Started and completed in March
    Transferred to finished goods in March
    Inventory in process, March 31
    Total

    Feedback
Answer #1

1) Ending work in process unit = 6000+108000-109500 = 4500 Units

2) Calculate equivalent unit

Whole Unit Direct Materials
Equivalent Units
Conversion Equivalent Units
Inventory in process, March 1 6000 0 6000*2/5 = 2400
Started and completed in March 103500 103500 103500
Transferred to finished goods in 109500 103500 105900
Inventory in process, March 31 4500 4500 4500*3/5 = 2700
Total 114000 108000 108600